Chocolate chip cookies are sold fo $6.95 per pound and white chocolate cookies for $5.95 per poud. How many pounds of chocolate chip cookies should be mixed with 4 pounds of white chocolate cookies to obtain a mixture that sells for $6.75 per pound?
Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

Chocolate chip cookies are sold fo $6.95 per pound and white chocolate cookies for $5.95 per poud. How many pounds of chocolate chip cookies should be mixed with 4 pounds of white chocolate cookies to obtain a mixture that sells for $6.75 per pound?
-------------------------
Value Equation:
value + value = value
6.95x + 5.95*4 = 6.75(x + 4)
Multiply thru by 100 to get:
695x + 595*4 = 675x + 675*4
20x = 80*4
x = 16 lbs (pounds of chocolate chip needed for the mixture)
==================
Cheers,
Stan H.

Answer by solver91311(24713)


First thing to do is change the dollars and cents to cents, eliminating the decimal fractions in your variable coefficients. It doesn't really matter in the end, it is simply neater and you are less prone to arithmetic error.

Let equal the required number of pounds of chocolate chip cookies.

The value of the chocolate chip cookies in the mix will then be

The value of the white chocolate cookies in the mix will be

The value of the mix has to be

But the value of the mix also has to be the sum of the values of the two ingredients, namely:



So set these two expressions equal to each other and solve for x.